Throwing a wrench. An angry construction worker throws
his wrench downward from a height of 128 feet with an
initial velocity of 32 feet per second. The height of the
wrench above the ground after t seconds is given
by S(t)=-16t^2- 32t+ 128.
a) What is the height of the wrench after 1 second?
b) How long does it take for the wrench to reach the
ground?

Answer by Alan3354(69443)

Throwing a wrench. An angry construction worker throws
his wrench downward from a height of 128 feet with an
initial velocity of 32 feet per second. The height of the
wrench above the ground after t seconds is given
by S(t)=-16t^2- 32t+ 128.
a) What is the height of the wrench after 1 second?
S(1) = -16 - 32 + 128
S(1) = 80 feet at 1 second
-----------------
b) How long does it take for the wrench to reach the
ground?
S(t) = -16t^2 - 32t + 128 = 0
t^2 + 2t - 8 = 0
(t+4)*(t-2) = 0
t = 2 seconds